I can't come up with a solution to this can some one help please?
X+Y+2Z=8
X+Y+5Z=11
X+6Y+4Z=30 THE TRIPLE SOLUTION SET IS?

Answer by scott8148(6628)

subtracting the 1st equation from the 2nd ___ 3Z = 3 ___ so, Z = 1

substituting gives ___ X + Y = 6
and ___ X + 6Y = 24

subtracting the two new equations gives ___ 5Y = 18 ___ so, Y = 3.6 (or 18/5)

substituting ___ X + 3.6 = 6 ___ X = 2.4 (or 12/5)